# Quantum Computing in Wireless Communications and Networking: A Tutorial-cum-Survey

### v1

@@ IV-C2 Applications of Conventional QC for Task Offloading
The previous works in this subsection overlook the mobility of end-uses or edge nodes even though in MEC or IoV scenarios. Given the rapid advancements in mobile communication technology, a diverse array of intelligent computation systems have surfaced, demanding more stringent QoS performance.
The channel resource allocation was presented in [193] in the background of task offloading in a three-tier, such as macro, pico and femto, MEC based heterogeneous 5G networks. The user equipment offloaded computation tasks to the local server, which significantly reduced the transmission latency. In the process of data transmission, multiple nodes competed for the channel resource. However, a random mobility model was designed to describe the movement of the user equipment. One approach in [192] introduced a joint offloading strategy based on QPSO for MEC-enabled vehicular networks. The strategy focused on minimizing delay costs and energy consumption by offloading tasks from vehicles to available service nodes, including service vehicles and nearby RSUs. To facilitate task offloading via V2V communication, a vehicle selection algorithm was introduced to determine an optimal offloading decision sequence. Moreover, to maintain population diversity in the QPSO algorithm, a crossover operator was employed to exchange information among individuals. However, it was assumed in the mobility model that vehicle speeds and coordinates were given in advance, presenting a practical limitation.
@@ IV-C3 Applications of AI-driven QC for Task Offloading
The emergence and rapid development of AI have transformed QC techniques for task offloading into a higher level of smartness and efficiency through many scenarios, e.g., IoV [194] and smart healthcare [195]. Specifically, in [194], a resource management strategy was introduced to achieve the balance between between communication and computation requirements by continuously monitoring the status of local vehicles, MEC-assisted WiFi links, and cloud-assisted 5G cellular networks. It empowered vehicles to make adaptive and flexible decisions regarding task offloading and transmission. To tackle the optimization problem in the face of time-varying channel conditions, it was reformulated as a MDP, solved by QRL to derive the real-time optimal resource management policy. The approach posited a virtual quantum server for performing optimization, whereas works [195, 196] explicitly introduced quantum servers.
@@
In [195], a dynamic computation offloading schedule was designed to minimize total latency within a hybrid setup, accommodating both classical (CPUs/GPUs) and quantum (QPUs) computing resources in cloud servers. This strategy ensured sustainable computation offloading while meeting individual latency constraints and the requisite success ratio for each task. Initially, the latency minimization problem was reformulated as a step-wise mixed-integer non-convex optimization task utilizing Lyapunov techniques. Subsequently, a DQN was employed for computation offloading mode selection, CPU/GPU mode or QPU mode, for cloud servers, addressing the optimization problem. Additionally, quantum edge servers were established in [196]. It formulated a classical and quantum computation offloading problem as a mixed-integer programming model, which was tackled using a multi-agent hybrid continuous-discrete DRL algorithm.
@@
QC coupled with DRL techniques also found extensive applications in addressing task offloading challenges across various scenarios, including IoT [197, 198], VR [199, 200], and Metaverse environments [201]. In [197], a MEC-enabled IoT framework was proposed, integrating QC capabilities. The architecture enabled IoT users to make sequential offloading and transmission decisions in an adaptive and flexible manner. The study formulated a joint optimal stochastic computation task offloading and dynamic resource allocation problem aimed at maximizing system energy efficiency. The optimization considered factors such as transmission latency, energy consumption, time-varying channel conditions, and data buffer constraints. Quantum state representations were employed to transform the continuous-time MDP problem into a QC optimization, which was effectively addressed by the proposed QDRL algorithm. The research team also extended their solution into 6G MEC based IoT system [198], but applying variational quantum circuits.
@@
Representative applications in 6G, VR and Metaverse, presented challenges in transmitting resources for large-volume streaming data, excessive computing resources for encoding and decoding operations, and extreme energy consumption in transmitting and receiving devices. QC coupled with DRL techniques emerged as a promising approach to address these challenges [199, 200, 201]. The study [199] presented an energy-aware resource management strategy for VR-supported industrial IoT. It addressed rendering task offloading and resource allocation challenges considering factors like content correlation, fluctuating channel conditions, and VR QoE. To handle this problem effectively, it was transformed into a sequential decision-making scenario under varying channel conditions over time. By defining a reward function that encapsulated energy consumption and VR QoE metrics, the optimization problem was formulated as a MDP, which was then tackled using the QRL algorithm to maximize the long-term system reward. The researchers extended their work to enhance resilience against malicious attacks by integrating the blockchain technique into the system, particularly in the context of VR-enabled medical treatment applications in 6G [200]. Additionally, they applied the QDRL algorithm instead of the QRL algorithm to enhance optimization efficiency. Another innovative application of 6G, Metaverse, leveraged DRL-driven QC algorithms to facilitate seamless immersion between reality and virtuality. Service subscribers accessed a real-time virtual world, constructed from images captured by on-site UAVs, within the Metaverse framework [201]. A key challenge arose in determining the optimal BSs to which UAVs offloaded their tasks to maximize benefits, particularly when BSs have limited bandwidth and computation resources. Specifically, to tackle the challenge posed by new environments or re-deployments, where traditional RL methods often encountered model collapse, collective learning was introduced. This approach enabled UAVs to leverage shared intelligence, facilitating rapid adaptation to new environments. The work can be improved by considering UAV mobility to achieve a more efficient system.
@@
In conclusion, QC offers valuable insights into task offloading, particularly in the context of evolving networking technologies like 6G, IoV, and UAV networks. Detailed summaries of these works are presented in Table X, with further elaboration on their QC algorithms available in Table XI. It is worth noting that task offloading within these scenarios typically involves distributed control and node mobility. However, only a few works provided solutions addressing these aspects, showed in Table XI. The primary reason for the first characteristic is current high overhead of deploying quantum machines. To the best of our knowledge, mobility remains an open challenge for task offloading integrated with QC.
@@ IV-D1 Overview
In the fast-paced world of digital content consumption, content caching, a technique widely employed by content delivery networks (CDNs), web servers, information centric network, etc., plays a pivotal role in optimizing the delivery of online content, profoundly influencing the overall user experience [202, 203]. It involves the strategic storage of frequently accessed content or computation results on servers distributed geographically, ensuring that users can access information with minimal latency, cost or other preferred objective. This practice significantly increases responsiveness and network bandwidth utilization.
Content caching scenarios are prevalent across various applications, spanning from edge computing [204], 5G cellular networks [205, 206], IoT [207] to video streaming [208, 209]. With the development of QC algorithms, the solutions leveraged QGA [204], QDL [210], and QRL [209] to tackle problems. The inherent technical challenges in content caching stem from several key factors. First, ensuring effective caching cooperation among different network entities is crucial for optimizing the distribution of cached content and maximizing overall network performance. Second, resource limitations, such as limited storage capacity and bandwidth availability, pose significant challenges in efficiently managing cached content and satisfying user demands. Third, the dynamic nature of network traffic and user preferences requires adaptive caching strategies that can adjust in real-time to changing conditions. Fourth, the heterogeneity of network devices and their capabilities introduces complexities in coordinating caching decisions across different devices with varying computational and storage resources. Finally, the distributed control manner of content caching systems necessitates robust algorithms and protocols for coordinating caching activities and ensuring seamless integration within the network architecture.
TABLE XII: Summary of QC Works on Content Caching in Wireless Networks
@@ IV-D2 Applications of Conventional QC for Content Caching
To alleviate backhaul link congestion and improve resource utilization in 5G ultra-dense cellular networks consisting of a macro, multiple picos and users, it is a traditional issue to cache content cooperatively in picos with resource limitations as illustrated in [204]. This study focused on an energy–delay tradeoff problem in collaborative edge computing-assisted and energy harvesting-powered ultra-dense networks. It regarded delay and energy as two types of cost and introduced a weighted cost function to transform the tradeoff problem into a cost minimization problem. An alternating optimization based on an improved QGA was proposed to solve the cost minimization problem. In the alternating optimization, the proposal divided the cost minimization problem into two sub-problems, adaptive tuning weight sub-problem and caching decision sub-problem. It improved QGA in terms of repairing unfeasible solutions and adaptively updating quantum genes. The problem arose from the fact the solutions derived from the individual sub-problems can not ensure the optimality of the original problem.
One of important features in 5G cellular networks is heterogeneity, i.e., small-cell base stations can be activated when necessary to be either a pico-cell base station or femto-cell base station [205], and cooperative caching performs in core gateways and BSs [206]. A joint caching and activation mechanism in [205] was designed to maximize system energy savings, while a cooperative caching policy was proposed in [206] to reduce the duplicate transmission and improve the green content delivery. Both two methods formulated the issue as an ILP problem, which were solved by QEA. Another heterogeneity came from assistance of UAVs for edge caching in 5G cellular networks, which had also received extensive attention from both industry and academia [208]. UAV-assisted edge caching is a promising technique to improve content transmission and enhance delivery efficiency by caching video contents closer to users. The study in [208] adaptively adjusted video quality based on time-varying network conditions and different user preferences according to proposed QEA. However, the work can be further enhanced by leveraging UAV mobility to dynamically adapt to user mobility patterns.
@@ IV-D3 Applications of AI-driven QC for Content Caching
AI-driven QC can significantly enhance the efficiency of content caching by leveraging its advantages in computing [210, 207, 209, 211]. In [210], DL was integrated into QC to address the traditional challenge of video caching with a detail process of training a QDL model. Specifically, a DL agent utilized the self-organizing maps algorithm to prioritize caching contents. The prioritized contents were stored in a quantum memory module using a two-level spin quantum phenomenon. Following the selection of the most suitable lattice map via self-organizing maps, data points below a specified threshold were mapped onto the data frame to retrieve the videos. These videos were considered high-priority for training based on the input data provided in the dataset. Similarly, the models for the other level prioritized content can be trained accordingly.
@@
Embedding RL into QC [207, 209] was another effective approach to solving complex problems in content caching, especially for the problems in sequential decision processes. In [207], there were two sub-problems of task offloading and content caching with limited resources in a traditional edge computing assisted network. The network architecture comprised a BS, edge servers, and IoT devices. The proposed approach aimed to optimize dynamic resource allocation, content caching strategy, and computation offloading policy to enhance energy efficiency in dynamic large-scale IoT networks. A QRL based algorithm was introduced to address the challenges posed by complex and dynamic environments, varying content popularity, and user preferences. However, there were two shortcomings in the proposal. First, the rationale behind jointly addressing the offloading and caching problem was unclear, lacking practical justification. Secondly, the solution relied on integer-based decisions for offloading and caching, potentially limiting its flexibility and scalability. QRL was also applied in the application of caching video streaming to improve QoE for users [209]. Its proposal jointly optimized content caching and spectrum allocation to minimize content delivery delay, taking into account time-varying content popularity, transmission method selection, and different requirements on EoE in a similar wireless network structure with [207]. The second shortcoming in [207] also arose in [209] from the binary caching and offloading decisions in its proposal. To address the same problem as [207], the study [211] proposed a solution by integrating DQN into QC, which involved the joint optimization of task offloading and content caching. However, a limitation arose as the two sub-problems were optimized independently, potentially leading to sub-optimal solutions for the original problem.
@@
In conclusion, QC algorithms are widely applied in the area of content caching, as presented in Table XII. There are still several issues in existing literature. Many related works implement their QC algorithms in computer simulations duo to the current developing stage of quantum computers. Additionally, current solutions often employ centralized control methods, primarily because computation services typically take the form of quantum clouds. Lastly, similar to the challenge in task offloading, node mobility remains an open issue when considering content caching.
@@ V-A Security and Privacy
Attacks on wireless pilot signals pose a significant threat to URLLC services. In grant-free URLLC scenarios, accurate allocation of distinct pilot resources to multiple users sharing the same time-frequency resource is crucial for the next-generation Node B to accurately identify these users under access collisions. This precise allocation is essential for maintaining accurate channel estimation, which is necessary for ensuring reliable data transmission. However, this procedure is susceptible to attacks targeting pilots. The research in [213] introduced an approach utilizing quantum learning to implement nonrandom superimposed coding for encoding and decoding pilots across multidimensional resources. The method enabled rapid learning and precise elimination of attack uncertainties. Specifically, pilots for uplink access by multiple users were encoded as distinct sub-carrier activation patterns. The Node B then decoded the desired pilots from observed sub-carrier activation patterns, which represented a superposition of sub-carrier activation patterns from accessing users. This was achieved through a joint design involving attack mode detection and user activity detection via a quantum learning network. The study in [214] solved the same security problem as [213] but in the area of random structure coding. It introduced an approach to encoding and decoding pilot signals using a random structured code across multidimensional physical resources. The proposed code employed random encoding with minimal structure to mitigate the impact of attacks. By leveraging group spatial channel features, the decoding process was modeled as a computational trap, leading to a security challenge akin to random computing with redundancy. To address this, the study utilized a quantum algorithm to learn the computational trap model, enabling the removal of computational redundancy and elimination of dispersed attacks. The proposal demonstrated the existence of a quantum black-box model corresponding to the computational trap and derived a precise expression for computational performance.
QC can also integrate with existing popular security and privacy techniques like blockchain [215] and federated learning [216, 217, 218] to enhance network performance. Inspired by classical federated learning techniques, the work in [216] aimed to enhance data privacy by aggregating updates from local computations to share model parameters. In pursuit of approximate optima in the parameter landscape, it introduced an extension of the traditional variational quantum algorithm. The variational quantum algorithms were also applied in future 6G space-air-ground integrated networks to address the complexity emerging from vast datasets and ML models aiming at data security and privacy [218].

### v3

@@ IV-C2 Applications of Conventional QC for Task Offloading
The previous works in this subsection overlook the mobility of end-uses or edge nodes even though in MEC or IoV scenarios. The channel resource allocation is presented in [172] in the background of task offloading in a three-tier, such as macro, pico and femto, MEC based heterogeneous 5G networks. User equipments offload computation tasks to the local server, which significantly reduces the transmission latency. In the process of data transmission, multiple nodes compete for the channel resource. However, a random mobility model is designed to describe the movement of the user equipment. One approach in [171] introduces a joint offloading strategy based on QPSO for MEC-enabled vehicular networks. The strategy focuses on minimizing delay costs and energy consumption by offloading tasks from vehicles to available service nodes, including service vehicles and nearby RSUs. To facilitate task offloading via V2V communication, a vehicle selection algorithm is introduced to determine an optimal offloading decision sequence. Moreover, to maintain population diversity in the QPSO algorithm, a crossover operator is employed to exchange information among individuals. However, it is assumed in the mobility model that vehicle speeds and coordinates are given in advance, presenting a practical limitation.
@@ IV-C3 Applications of AI-driven QC for Task Offloading
The emergence and rapid development of AI have transformed QC techniques for task offloading into a higher level of smartness and efficiency through many scenarios, e.g., IoV [173] and smart healthcare [174]. Specifically, in [173], a resource management strategy is introduced to achieve the balance between communication and computation requirements by continuously monitoring the status of local vehicles, MEC-assisted WiFi links, and cloud-assisted 5G cellular networks. It empowers vehicles to make adaptive and flexible decisions regarding task offloading and transmission. To tackle the optimization problem in the face of time-varying channel conditions, the problem is reformulated as a MDP problem, solved by QRL to derive the real-time optimal resource management policy. The approach assumes a virtual quantum server for performing optimization, whereas works [174, 175] explicitly introduce quantum servers.
@@
In [174], a dynamic computation offloading schedule is designed to minimize total latency within a hybrid setup, accommodating both classical (CPUs/GPUs) and quantum (QPUs) computing resources in cloud servers. This strategy ensures sustainable computation offloading while meeting individual latency constraints and the requisite success ratio for each task. Initially, the latency minimization problem is reformulated as a step-wise mixed-integer non-convex optimization task by utilizing Lyapunov techniques. Subsequently, DQN is employed for computation offloading mode selection, CPU/GPU mode or QPU mode, for cloud servers. Additionally, quantum edge servers are established in [175]. It formulates a classical and quantum computation offloading problem as a mixed-integer programming model, which is tackled using a multi-agent hybrid continuous-discrete DRL algorithm.
@@
QC coupled with DRL techniques also finds extensive applications in addressing task offloading challenges across various scenarios, including IoT [176, 177], VR [178, 179], and Metaverse environments [180]. In [176], a MEC-enabled IoT framework is proposed, integrating QC capabilities. The architecture enables IoT users to make sequential offloading and transmission decisions in an adaptive and flexible manner. The study formulates a joint optimal stochastic computation task offloading and dynamic resource allocation problem aimed at maximizing system energy efficiency. The optimization considers factors such as transmission latency, energy consumption, time-varying channel conditions, and data buffer constraints. However, in highly dynamic and complex IoT environments, a Q-table in classical DRL algorithms face significant challenges. The stochastic and time-varying nature of channel conditions, combined with a large search space, leads to high time and space complexities. The discretization of continuous state and action spaces in the Q-table further hinders learning efficiency, primarily due to the high dimensionality. Additionally, the slow learning speed and limited storage capacity of classical DRL models are inadequate to meet the stringent offloading decision-making and processing demands of IoT systems. Quantum state representations are employed to transform the continuous-time MDP problem into QC optimization, which is effectively addressed by the proposed QDRL algorithm. All potential quantum representation states can be modified in magnitude through a set of orthonormal eigenstates. Leveraging quantum parallelism, a unitary matrix operates on the superposition of mixed qubits, simultaneously affecting all states. The research team also extends their solution into 6G MEC based IoT system [177], but applying variational quantum circuits.
@@
Representative applications in 6G, VR and Metaverse, presented challenges in transmitting resources for large-volume streaming data, excessive computing resources for encoding and decoding operations, and extreme energy consumption in transmitting and receiving devices. QC coupled with DRL techniques emerges as a promising approach to address these challenges [178, 179, 180]. The study [178] presents an energy-aware resource management strategy for VR-supported industrial IoT. It addresses rendering task offloading and resource allocation challenges considering factors like content correlation, fluctuating channel conditions, and VR QoE. To handle this problem effectively, the problem is transformed into a sequential decision-making scenario under varying channel conditions over time. By defining a reward function that encapsulates energy consumption and VR QoE, the optimization problem is formulated as an MDP, which is then tackled using QRL to maximize the long-term system reward. The researchers extend their work to enhance resilience against malicious attacks by integrating the blockchain technique into the system, particularly in the context of VR-enabled medical treatment applications in 6G [179]. Additionally, they apply QDRL instead of QRL to enhance optimization efficiency. Similar to the work [176], the study transforms the joint resource allocation based on RL into QDRL to overcome the challenge of ”curse of dimensionality”. It chooses an $n$-qubit quantum system and uses $n$ qubits to represent the eigen actions. An entangled quantum action is constructed by an orthogonal set of base vectors that are called eigen actions. Each eigen action corresponds to a classic action in RL. Then, the classic actions can be transformed into an entangled quantum action using the base vectors. The states and actions in RL is transformed into a quantum-collapse-based mapping according to the probability amplitude. Finally, the quantum searching for the optimal actions is performed according to the Grover iteration based on Hadamard gate.
Another innovative application of 6G, Metaverse, leverages DRL-driven QC algorithms to facilitate seamless immersion between reality and virtuality. Service subscribers access a real-time virtual world, constructed from images captured by on-site UAVs, within the Metaverse framework [180]. A key challenge arises in determining the optimal BSs to which UAVs offloaded their tasks to maximize benefits, particularly when BSs have limited bandwidth and computation resources. Specifically, to tackle the challenge posed by new environments or re-deployments, where traditional RL methods often encountered model collapse, collective learning is introduced. This approach enables UAVs to leverage shared intelligence, facilitating rapid adaptation to new environments. The work can be improved by considering UAV mobility to achieve a more efficient system.
@@
In conclusion, QC offers valuable insights into task offloading, particularly in the context of evolving networking technologies like 6G, IoV, and UAV networks. Detailed summaries of these works are presented in Table XI, with further elaboration on their QC algorithms available in Table XII. It is worth noting that task offloading within these scenarios typically involves distributed control and node mobility. However, only a few works provide solutions to address these aspects, showed in Table XII. The primary reason for the first characteristic is current high overhead of deploying quantum machines. To the best of our knowledge, mobility remains an open challenge for task offloading integrated with QC.
@@ IV-D1 Overview
In the fast-paced world of digital content consumption, content caching, a technique widely employed by content delivery networks (CDNs), web servers, information centric network, etc., plays a pivotal role in optimizing the delivery of online content, profoundly influencing the overall user experience [181]. It involves the strategic storage of frequently accessed content or computation results on servers distributed geographically, ensuring that users can access information with minimal latency, cost or other preferred objective. This practice significantly increases responsiveness and network bandwidth utilization.
Content caching scenarios are prevalent across various applications, spanning from edge computing [182], 5G cellular networks [183], IoT [184] to video streaming [185]. With the development of QC algorithms, the solutions leverage QGA [182], QDL [186], and QRL [187] to tackle problems. The inherent technical challenges in content caching stem from several key factors. First, ensuring effective caching cooperation among different network entities is crucial for optimizing the distribution of cached content and maximizing overall network performance. Second, resource limitations, such as limited storage capacity and bandwidth availability, pose significant challenges in efficiently managing cached content and satisfying user demands. Third, the dynamic nature of network traffic and user preferences requires adaptive caching strategies that can adjust in real-time to changing conditions. Finally, the distributed control manner of content caching systems necessitates robust algorithms and protocols for coordinating caching activities and ensuring seamless integration within the network architecture.
TABLE XIII: Summary of QC Works on Content Caching in Wireless Networks
@@ IV-D2 Applications of Conventional QC for Content Caching
To alleviate backhaul link congestion and improve resource utilization in 5G ultra-dense cellular networks consisting of a macro, multiple picos and users, it is a traditional issue to cache content cooperatively in picos with resource limitations as illustrated in [182]. This study focuses on an energy–delay tradeoff problem in collaborative edge computing-assisted and energy harvesting-powered ultra-dense networks. It regards delay and energy as two types of cost and introduces a weighted cost function to transform the tradeoff problem into a cost minimization problem. An alternating optimization based on an improved QGA is proposed to solve the cost minimization problem. The proposal divides the cost minimization problem into two sub-problems, adaptive tuning weight sub-problem and caching decision sub-problem. It improves QGA in terms of repairing unfeasible solutions and adaptively updating quantum genes. The problem arises from the fact the solutions derived from the individual sub-problems can not ensure the optimality of the original problem.
One of important features in 5G cellular networks is heterogeneity, i.e., small-cell BSs can be activated when necessary to be either a pico-cell BS or femto-cell BS [188], and cooperative caching performs in core gateways [183]. A joint caching and activation mechanism in [188] is designed to maximize system energy savings, while a cooperative caching policy is proposed in [183] to reduce the duplicate transmission and improve the green content delivery. Both two methods formulate the issue as an ILP problem, which is solved by QEA. Another heterogeneity comes from assistance of UAVs for edge caching in 5G cellular networks, which has also received extensive attention from both industry and academia [185]. UAV-assisted edge caching is a promising technique to improve content transmission and enhance delivery efficiency by caching video closer to users. The study [185] adaptively adjusts video quality based on time-varying network conditions and different user preferences according to proposed QEA. However, the work can be further enhanced by leveraging UAV mobility to dynamically adapt to user mobility.
@@ IV-D3 Applications of AI-driven QC for Content Caching
AI-driven QC can significantly enhance the efficiency of content caching by leveraging its advantages in computing [186, 184, 187, 189]. In [186], DRL is integrated into QC to address the traditional challenge of video caching with a detail process of training a QDL model. Specifically, a DRL agent utilizes the self-organizing map algorithm to prioritize caching contents. The prioritized contents are stored in a quantum memory module using a two-level spin quantum phenomenon. Following the selection of the most suitable lattice map via self-organizing maps, data points below a specified threshold are mapped onto the data frame to retrieve the videos. These videos are considered high-priority for training based on the input data provided in the dataset. Similarly, the models for the other level prioritized content can be trained accordingly.
@@
Embedding RL into QC [184, 187] is another effective approach to solving complex problems in content caching, especially for the problems in sequential decision processes. In [184], there are two sub-problems of task offloading and content caching with limited resources in a traditional edge computing assisted network. The network architecture comprises a BS, edge servers, and IoT devices. The proposed approach aims to optimize dynamic resource allocation, content caching strategy, and computation offloading policy to enhance energy efficiency in dynamic large-scale IoT networks. A QRL based algorithm is introduced to address the challenges posed by complex and dynamic environments, varying content popularity, and user preferences. However, there are two shortcomings in the proposal. First, the rationale behind jointly addressing the offloading and caching problem is unclear, lacking practical justification. Secondly, the solution relies on integer-based decisions for offloading and caching, potentially limiting its flexibility and scalability. QRL is also applied in the application of caching video streaming to improve QoE [187]. Its proposal jointly optimizes content caching and spectrum allocation to minimize content delivery delay, taking into account time-varying content popularity, transmission method selection, and different requirements on QoE in a similar wireless network structure with [184]. The second shortcoming in [184] also arises from the binary caching and offloading decisions in its proposal. To address the same problem as [184], the study [189] proposes a solution by integrating DQN into QC, which involves the joint optimization of task offloading and content caching. However, a limitation arises as the two sub-problems are optimized independently, potentially leading to sub-optimal solutions for the original problem.
@@
In conclusion, QC algorithms are widely applied in the area of content caching, as presented in Table XIII. There are still several issues in existing literature. Many related works implement their QC algorithms in computer simulations duo to the current developing stage of quantum computers. Additionally, current solutions often employ centralized control methods, primarily because computation services typically take the form of quantum clouds. Lastly, similar to the challenge in task offloading, node mobility remains an open issue when considering content caching.
@@ V-A Security and Privacy
Attacks on wireless pilot signals pose a significant threat to URLLC services. In grant-free URLLC scenarios, accurate allocation of distinct pilot resources to multiple users sharing the same time-frequency resource is crucial for the next-generation eNodeB to accurately identify these users under access collisions. This precise allocation is essential for maintaining accurate channel estimation, which is necessary for ensuring reliable data transmission. However, this procedure is susceptible to attacks targeting pilots. The research in [191] introduces an approach utilizing quantum learning to implement nonrandom superimposed coding for encoding and decoding pilots across multidimensional resources. The method enables rapid learning and precise elimination of attack uncertainties. Specifically, pilots for uplink access by multiple users are encoded as distinct sub-carrier activation patterns. The enodeB then decodes the desired pilots from observed sub-carrier activation patterns that represent a superposition. This is achieved through a joint design involving attack mode detection and user activity detection via a quantum learning network. The study in [192] solves the same security problem as [191] but in the area of random structure coding. It introduces an approach to encoding and decoding pilot signals using a random structured code across multidimensional physical resources. The proposed code employs random encoding with minimal structure to mitigate the impact of attacks. By leveraging group spatial channel features, the decoding process is modeled as a computational trap, leading to a security challenge parallel with random computing with redundancy. To address this, the study utilizes a quantum algorithm to learn the computational trap model, enabling the removal of computational redundancy and elimination of dispersed attacks. The proposal demonstrates the existence of a quantum black-box model corresponding to the computational trap and derives a precise expression for computational performance.
QC can also integrate with existing popular security and privacy techniques like blockchain [193] and federated learning [194, 195, 196] to enhance network performance. Inspired by classical federated learning techniques, the work in [194] aims to enhance data privacy by aggregating updates from local computations to share model parameters. To obtain approximate optima in the parameter landscape, it introduces an extension of the traditional variational quantum algorithm. The variational quantum algorithms are also applied in future 6G space-air-ground integrated networks to address the complexity emerging from vast datasets and ML models aiming at data security and privacy [196].